1. Above 882 oC, titanium has a BCC crystal structure, with a = 0.332 nm. Below this temperature, titanium has a HCP structure, with a = 0.2978 nm and c = 0.4735 nm. Determine the percent volume change when BCC titanium transforms to HCP titanium. Is this a contraction or expansion?
2. The element lithium has BCC packing with a body-centered cubic unit cell. The volume of the unit cell is 4.32 x 10−26 L. Calculate the density (g/cm3) of the element.
3. What is the volume of the unit cell? A metal crystallizes in a face-centered cubic lattice. The radius of the atom is 0.197 nm. The density of the element is 1.54 g/cm3.
4. An element crystallizes in a body centered cubic unit cell. The radius of this element is 2.2 x 10−10 m, and the density is 0.991 g/cm3. How many moles of atoms are within one unit cell?
5. Nickel has a FCC crystal structure, an atomic radius of 0.125 nm, and an atomic weight of 58.68 g/mol. Compute and compare its theoretical density with the experimental value of 8.90 g/cm3.
1
Expert's answer
2020-12-01T09:29:01-0500

Q146799


1. Above 882 oC, titanium has a BCC crystal structure, with a = 0.332 nm. Below this temperature, titanium has a HCP structure, with a = 0.2978 nm and c = 0.4735 nm. Determine the percent volume change when BCC titanium transforms to HCP titanium. Is this a contraction or expansion?


Solution :


Step 1 : Volume of 1 unit of BCC


Volume per unit cell of BCC crystal is given by, V BCC = a3 ;


We are given a = 0.332 nm. Substitute this in the formula, we have


VBCC = (0.332 nm) 3 = 0.036594 nm3 ;



Step 2 : Volume of 1 unit of HCP

We are given a = 0.2978 nm and c = 0.4735 nm


Volume per unit cell of HCP is given by,


"V_{HCP} = \frac{3\sqrt{3}*a^2*c}{2}"



"V_{HCP} = \frac{3\sqrt{3}*0.2978^2*0.4735}{2}"



VHCP = 0.1091 nm2 ;  


Step 3 : To find volume per unit atom for both BCC and HCP ;


Each unit of BCC contains 2 atoms and each units of HCP contains 6 atoms.  


Volume per unit atom for BCC = 0.036594 nm3 /2= 0.01830 nm3 / atom.  


volume per unit atom for HCP = 0.1091 nm2 /6= 0.018183 nm3 / atom ;  


Step 4 : To find the % change in volume due to allotropic transformation of Titanium  


% change in volume = "\frac{V_{HCP}-V_{BCC}}{V_{BCC}}*100;"


= "\frac{0.018183nm^3-0.01830nm^3}{0.01830nm^3}*100" ;


= -0.639 % ;



Hence the % change in volume when BCC titanium transforms to HCP titanium = - 0.639 % .


The negative sign implies that there is a contraction when Titanium undergoes allotropic transformation from BCC to HCP.



2. The element lithium has BCC packing with a body-centered cubic unit cell. The volume of the unit cell is 4.32 x 10−26 L. Calculate the density (g/cm3) of the element.
Solution :


Each unit of BCC contains 2 atoms. So there would be 2 atoms of lithium in each unit cell.


Atomic mass of Li = 6.941g/mol ;


Step 1 : To find mass of single unit of BCC


Using Avagadro’s number we will find mass of 1 atom of Li.


Mass of 1 atom of Li = 6.941g Li /1mol Li * 1 mol Li /6.022 * 1023 atoms of Li  

= 1.1526 * 10-23 grams/ atom .  


Mass of single BCC unit of Li = 2 * mass of 1 atom of Li = 2 * 1.1526 * 10-23 grams/ atom


= 2.3052 * 10-23 grams.



Step 2 : Convert the given volume to cm3 .

In question, we are given the volume of single unit of BCC = 4.32 x 10−26 L


Convert this to cm3 , using 1 L = 1000 cm3 ;


volume of single unit in cm3 = 4.32 x 10−26 L * 1000 cm3 /1L


= 4.32 x 10−23 cm3 ;


Step 3: Find density using mass and volume.  


"density = \frac{mass}{volume};"


"density = \frac{2.3052 * 10^{-23} grams}{ 4.32 x 10^{−23} cm^{3}}"



density = 0.5336 g/cm3 ;  


In question, we are given volume in 3 significant figure.

So our final answer must also be in 3 significant figure.

Hence the density of lithium is 0.534 g/cm3



3. What is the volume of the unit cell? A metal crystallizes in a face-centered cubic lattice. The radius of the atom is 0.197 nm. The density of the element is 1.54 g/cm3.

Solution:


For FCC unit cell,  


"a=2R\sqrt{2}"

"V_{FCC}=a^3 = (2R\sqrt{2} )^3"


"V_{FCC}=a=(2*0.197*\sqrt{2})^3 ;"


"V_{FCC}=0.173 nm^3;"


So the volume of the unit cell is 0.173 nm3 ;



4. An element crystallizes in a body centered cubic unit cell. The radius of this element is 2.2 x 10−10 m, and the density is 0.991 g/cm3. How many moles of atoms are within one unit cell?

Solution :


Each unit cell of Body centered cubic system contains 2 atoms.


and we know, 1 mol = 6.022 * 1023 atoms.


So, moles of atoms in a single unit cell = 2 atoms * 1 mol /6.022 * 1023 atoms.

= 3.32 * 10-24 moles.


So there are 3.3 * 10-24 moles of atoms in one unit cell of BCC.



5. Nickel has a FCC crystal structure, an atomic radius of 0.125 nm, and an atomic weight of 58.68 g/mol. Compute and compare its theoretical density with the experimental value of 8.90 g/cm3.


Solution :


Step 1: Find the volume of single unit of FCC.


For FCC crystal structure,


"V_{FCC}=a^3 = (2*R*\sqrt{2})^{3}" ;


Convert 0.125 nm to cm;


Radius of atom in 'cm' = 0.125nm * 1m/10-9 nm * 100cm/1m = 0.125 *10-7 cm;


"V_{FCC}=a^3 = (2*0.125*10^{-7}*\sqrt{2})^{3};"


"V_{FCC} = (3.5355*10^{-8})^3;"


"V_{FCC}= 4.419 *10^{-23};"


Step 2 : To find mass of 1 unit of FCC.


Each unit of FCC contains 4 atoms.


Mass of 1 atom = 58.68 g/1mol * 1 mol/ 6.022 * 1023 atoms = 9.744 * 10-23 g/ atom.



So, mass of 1 unit of FCC = 4 * 9.744 * 10-23 g/ atom = 3.898 * 10-22 grams.


Step 3 : To find the theoretical density.  


"density=\frac{mass}{volume} = \frac{3.898*10^{-22}g}{4.419*10^{-23}cm^{3}}"



Theoretical density = 8.819 g/cm3 ;

which in 3 significant figure is 8.82 g/cm3 ;


The experimental density given in the question is 8.90 g/cm3 ;


The Theoretical density is less than the experimental density.